OpenStudy (anonymous):

What is 3 1/4 plus 3 3/8? Medal and fan!

OpenStudy (anonymous):

First, what is 3 1/4?

OpenStudy (anonymous):

a fraction

OpenStudy (anonymous):

haha true, but what is the value?

OpenStudy (anonymous):

hint: 1=4/4

OpenStudy (anonymous):

three is the number and one fourths is equal to 0.25

OpenStudy (anonymous):

3 1/4 = 13/4

OpenStudy (anonymous):

do you understand that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, now do the same thing for 3 3/8

OpenStudy (anonymous):

okay I think 27/8

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):